-2(5x + 4)/ 3 = -3(3x + 2) - 7/3
Help!

Answer by rfer(16322)

(-10x-8)/3=(-9x-6)-7/3
-10x-8=((-9x-6)-7/3)(3)
-10x-8=-27x-18-7
-10x+27x=-18-7+8
17x=-17
x=-17/17
x=-1